# Faster Reinforcement Learning by Freezing Slow States

### v2

@@ Introduction
See Appendix D.2. ∎
This result above can be interpreted as the regret being bounded by
@@
The main result of this section follows directly from Lemma 6.1 and is given in Corollary 6.1, which shows the expected regret $\mathcal{R}(\tilde{\mu}^{*},\tilde{\bm{\pi}}^{*})$ of applying the policy learned from the frozen-state hierarchical approximation to the base model.
Corollary 6.1.
Let $\langle\mathcal{S},\mathcal{A},\mathcal{W},f,r,\gamma\rangle$ be a $(\alpha,d_{\mathcal{Y}})$-fast-slow MDP satisfying Assumption 2. The regret of applying the $T$-periodic policy $(\tilde{\mu}^{*},\tilde{\bm{\pi}}^{*})$ in the base model is bounded by
$\mathcal{R}(\tilde{\mu}^{*},\tilde{\bm{\pi}}^{*})\leq\biggl{(}\frac{2\gamma^{T}}{(1-\gamma^{T})^{2}}+\frac{2}{1-\gamma^{T}}\biggr{)}\epsilon_{r}(\gamma,\alpha,d_{\mathcal{Y}},\mathbf{L},T)+\biggl{(}\frac{2\gamma^{2T}}{(1-\gamma^{T})^{2}}+\frac{2\gamma^{T}}{1-\gamma^{T}}\biggr{)}L_{U}\,d(\alpha,d_{\mathcal{Y}},T)$
where $d(\alpha,d_{\mathcal{Y}},T)=2d_{\mathcal{Y}}(\alpha+1)(T-1)$.
Proof.
We apply Lemma 6.1 with $\bm{\pi}=\tilde{\bm{\pi}}^{*}$, $J_{1}=J_{1}^{*}$, and $V=V^{*}(J_{1}^{*},\tilde{\bm{\pi}}^{*})$, while noting that by Proposition 6.1, $\epsilon_{r}(\bm{\pi}^{*},J_{1})\leq\epsilon_{r}(\gamma,\alpha,d_{\mathcal{Y}},\mathbf{L},T)$. ∎
Regret of Base VI
Proposition 6.2 is a well-known property that gives the required number of iterations of exact VI on the base model needed for the resulting policy to achieve a desired level of regret.
Proposition 6.2.
we recall that $r_{\textnormal{max}}$ is an upper bound on the reward.
Proof.
See Appendix E.2. ∎
@@
An instance of FSVI is associated with two primary quantities: $k$, the number of VI iterations, and $T$, the number of periods the slow state is frozen in the frozen-state approximation. The next theorem makes use of this lemma to provide a bound on the regret of the policy obtained for a particular $k$ and $T$.
@@
Regardless, such regret analysis is valuable, as it can help us understand and clarify the trade-offs involved when selecting various parameters, primarily $T$, the horizon of the frozen lower-level problem. This is particularly true in our setting, where we are balancing several sources of error (reward error and end-of-horizon error) with computational improvements. In this section, we numerically compute the regret bounds for various values of $T$ in order to obtain insights.
First, we examine the reward approximation error analyzed in Proposition 6.1. Depending on how the “freeze error” and the “end-of-horizon error” interact, we see different behaviors of the overall reward approximation error, as illustrated in Figure 4. For both examples, the end-of-horizon error increases at first, but then decreases toward zero as $T$ becomes large. In the example with $L_{f}<1$ (left), the freeze error converges as $T$ increases; however, if $L_{f}>1$, the freeze error grows without bound. This leads to two distinct patterns in the overall reward approximation error: (1) when the freeze error is well-controlled, the total reward approximation error generally follows the unimodal shape of the end-of-horizon error; and (2) when the freeze error grows without bound, the reward approximation error follows a unimodal shape at first, but eventually grows again as the freeze error accumulates.
Next, we show how Proposition 6.2 (the regret of Base VI) and Theorem 6.1 (the regret of FSVI) can be combined to provide intuition about whether to use Base VI or FSVI (and if so, which $T$?). We plot regret versus computational cost, using the definitions in Section 5.2. For each regret result, we plot the regret versus computational cost as the number of value iteration steps $k$ increases ($k$ is not shown). This exercise leads us to interesting takeaways illustrated in Figure 5:
@@
1
Input: A distribution over states $\rho$, number of iterations $k$.
@@ 4       Sample a dataset $\mathcal{D}=\{s^{(i)}\}_{i=1}^{N}\sim\rho$
6             $y^{(i)}=\max_{a}r(s^{(i)},a)+\gamma\,\mathbb{E}\bigl{[}U^{i-1}(f(s^{(i)},a,w))\bigr{]}$
@@ 8       $U^{i}=\texttt{Regress}(\{(s^{(j)},y^{(j)})\}_{j=1}^{N})$
11       $\nu^{k}(s)=\operatorname*{arg\,max}_{a}r(s,a)+\gamma\,\mathbb{E}\bigl{[}U^{k}(f(s,a,w))\bigr{]}$
@@
13
Algorithm 3 Fitted Value Iteration (Munos and Szepesvári, 2008)
1
Input: State distribution $\rho$, number of iterations $k$, number of basepoints $N$, number of next-state samples $M$.
@@
2 Initialize $\hat{J}_{T}\equiv 0$.
3
4for $t=T-1,T-2,\ldots,1$ do
5       Sample basepoints $\mathcal{D}_{t}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
6       for each $(x^{(i)},y^{(i)})\in\mathcal{D}_{t}$ do
7             $J_{t}^{(i)}=\max_{a\in\mathcal{A}}\frac{1}{M}\sum_{j=1}^{M}\left[r^{(i,a,j)}+\gamma\,\tilde{J}_{t+1}\left(x^{(i)},y^{(i,a,j)}\right)\right]$ where $y^{(i,a,j)}$ and $r^{(i,a,j)}$ are sampled from the generative model, conditional on $x^{(i)}$ and action $a$.
@@ 5       Sample basepoints $\mathcal{D}_{t}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
9       $\tilde{J}_{t}=\texttt{Regress}\Bigl{(}\bigl{\{}\bigl{(}(x^{(i)},y^{(i)}),\tilde{J}_{t}^{(i,j)}\bigr{)}\bigr{\}}_{i=1}^{N}\Bigr{)}$
@@
11Let $\tilde{\bm{\pi}}^{*}$ be greedy with respect to $\tilde{J}_{t}$, for each $t$.
12 Initialize $\hat{V}^{0}$ arbitrarily. for $k^{\prime}=1,2,\ldots,k$ do
13       Sample basepoints $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
@@
15             ${V}^{(k^{\prime})}=\max_{a\in\mathcal{A}}\frac{1}{M}\sum_{j=1}^{M}\left[\tilde{R}^{(i,a,j)}+\gamma^{T}{V}^{k^{\prime}-1}\left(x^{(i,a,j)},y^{(i,a,j)}\right)\right]$ where $x^{(i,a,j)}$, $y^{(i,a,j)}$, and $\tilde{R}^{(i,a,j)}$ are sampled using a $T$-step transition conditioned on $x^{(i)}$, $a$, $\tilde{\bm{\pi}}^{*}$ and $\tilde{J}_{1}$.
@@
19for $s_{0}=(x_{0},y_{0})\in\mathcal{S}$ do
20       $\tilde{\mu}^{k}(x_{0},y_{0})=\operatorname*{arg\,max}_{a\in\mathcal{A}}\frac{1}{M}\sum_{j=1}^{M}\left[\tilde{R}^{(i,a,j)}+\gamma^{T}\tilde{V}^{k}\left(x_{T}^{(i,a,j)},y_{T}^{(i,a,j)}\right)\right]$ where $x_{T}^{(i,a,j)}$, $y_{T}^{(i,a,j)}$, and $\tilde{R}^{(i,a,j)}$ are sampled from a $T$-step transition as above.
@@
So far, we have operated in the dynamic programming setting, where the model (rewards and transitions) is assumed to be known. We now move on to a more practical and realistic setting where we assume access to a generative model or simulator of the environment (also referred to as a restart distribution), an assumption frequently used in the reinforcement learning (RL) literature (Kakade and Langford, 2002; Kearns et al., 2002; Munos and Szepesvári, 2008; Li et al., 2020; Agarwal et al., 2020). This setting is widely studied in the RL literature and captures many real-world use cases where it is possible to simulate environment transitions (Bellemare et al., 2012; Todorov et al., 2012; Brockman et al., 2016; Tassa et al., 2018) but not to access closed-form model dynamics. This is an important setting that falls in-between dynamic programming and traditional online RL.
Fitted value iteration (FVI) is a simple method that approximates the value iteration backup step using regression (Munos and Szepesvári, 2008), which we keep as an abstract subroutine using the notation $\texttt{Regress}\bigl{(}\{(s^{(j)},y^{(j)})\}_{j=1}^{N}\bigr{)}$, which refers to the problem of regressing $\{y^{(j)}\}$ onto $\{s^{(j)}\}$. In Algorithm 3, we show the abstract version of FVI and in Algorithm 4, we give the frozen-state extension, Frozen-state FVI (FSFVI), which can be viewed as a practical extension of FSVI for the RL setting.
Additionally, in Appendix B, we present a version of FVI that is less practical, but more amenable to theoretical analysis, leading to a regret bound that we provide in Theorem B.1. These results are motivated by the approximate value iteration method first presented in Tsitsiklis and Van Roy (1996). Given that both the method and analysis are technical and notation-heavy, we relegated the discussion to the Appendix.
@@
For our numerical experiments, we operate in the practical simulated-based RL setting described above, where a generative model or simulator is available but the MDP model is not explicitly known. In Section 5.2, we described the computational cost of FSVI for the setting with a known model (with expectations computed exactly). However, for the RL setting, we opt to use the number of value function evaluations as our primary proxy for computational effort. This is motivated by the observation that in any value-based RL method—whether value- or Q-iteration, or their “fitted” variants—each Bellman update involves evaluating a previous iteration’s value function on samples from the simulator. We compare to a range of algorithms. For purely tabular methods, we consider:
  • FSVI with $T\in\{3,6,12\}$: Frozen-state value iteration as introduced Algorithm 2. For each domain, we run FSVI for each $T$ value in order to illustrate the trade offs.
  • Base VI: Exact value iteration as introduced in Algorithm 1. We also use Base VI to compute an approximation of the optimal value.
  • QI: Q-iteration is the analogue of VI using a Q-function representation.
  • Slow-agnostic VI: A simple baseline that pretends the slow state does not exist and assumes the value function only depends on the fast state. [^7] To implement this method, we convert the “full” state to a “partial” state by deleting the slow state, and when needing to interact with the simulator, we convert from a “partial” state back to a “full” state by injecting a random value for the slow state.
@@
  • {Lin, Deep}-FSFVI: Our proposed frozen-state fitted value iteration, as introduced in Algorithm 4, using either linear function approximator (Lin-FSFVI) or a neural network function approximator (Deep-FSVI).
  • {Lin, Deep}-FVI: Fitted value iteration (Munos and Szepesvári, 2008) using either linear function approximator (Lin-FVI) or a neural network function approximator (Deep-FVI). This algorithm is also known as least-square Monte Carlo, popularized by Longstaff and Schwartz (2001) and frequently used in the operations literature (see, e.g., Nadarajah et al. (2017)).
  • {Lin, Deep}-FQI: Fitted Q-iteration (Ernst et al., 2005) using either linear function approximator (Lin-FQI) or a two-layer neural network function approximator (Deep-FQI).
We test these algorithms across three domains, summarized in Table 2. Each domain uses a high discount factor of $\gamma=0.995$, corresponding to an effective planning horizon of approximately $2001/(1-\gamma)=$ 200 steps. We describe each experiment below; in each section, we use self-contained notation. Detailed experimental settings are given in Appendix A.
@@
Inventory control 561 11 6,171 Base VI, FSVI, QI, Slow-agnostic VI
Spatial-task gridworld 4,356 32 139,392 Base VI, FSVI, QI, Slow-agnostic VI
@@
where $d_{t+1}$ is drawn from a distribution that depends on $d_{t}$. The reward in each period is given by:
@@
which represents the revenue, variable order cost, and fixed order cost. The objective is to maximize infinite horizon discounted reward, with $\gamma=0.995$. In our experiment, we set $y_{\textnormal{max}}=50$, $\mathcal{A}_{\textnormal{order}}=\{0,5,\ldots,50\}$, $d_{t+1}=\textnormal{clip}_{[0,50]}(d_{t}+\epsilon_{t+1})$, and $\{\epsilon_{t}\}$ are i.i.d. random variables that take value 0 with probability $0.8$ and ±$\pm 1$ with probability 0.1 each. [^8] We define $\textnormal{clip}_{[a,b]}(x)=\max\{\min\{x,b\},a\}$.
Figure 6: Mean return (with 95% CI) versus computational cost for the inventory problem across five independent replications.
The results, shown in Figure 6, demonstrate that frozen-state methods substantially outperform standard approaches in terms of computational efficiency. Specifically, FSVI with moderate values of $T$, such as $T=6$ and $T=12$, achieves near-optimal return with significantly fewer value function evaluations compared to Base VI and QI. This supports the intuition that freezing the slow-moving demand state $d_{t}$ over short planning horizons can preserve essential structure while simplifying the lower-level planning problem.
Among the FSVI variants, we find that $T=6$ offers the best overall performance when considering both reward and computational cost. Larger values of $T$, such as $T=12$, initially converge faster but eventually suffer from increased approximation error due to longer freezing horizons. Conversely, if we look at the earlier part of the curves, FSVI with $T=3$ seems to benefit less from the improved discount factor.
While Base VI and QI eventually reach optimal performance, they require significantly more computation. In contrast, the slow-agnostic baseline performs poorly, confirming that ignoring the slow state leads suboptimal policies. These findings reinforce both our theoretical claims from Section 6 and the trade offs we observed in Figure 5 from Section 7.

### v3

@@ A General Lemma
See Appendix E.3. ∎
The result of Lemma 6.1 above can be interpreted as the regret being bounded by
@@
An instance of FSVI is associated with two primary quantities: $k$, the number of VI iterations, and $T$, the number of periods the slow state is frozen in the frozen-state approximation. The next theorem gives a bound on the regret of the policy obtained for a particular $k$ and $T$.
@@
Regardless, such regret analysis is valuable, as it can help us understand and clarify the trade-offs involved when selecting various parameters, primarily $T$, the horizon of the frozen lower-level problem. This is particularly true in our setting, where we are balancing several sources of error with computational improvements. In this section, we numerically compute the regret bounds for various values of $T$ in order to obtain insights.
Illustrating Reward Error
First, we examine the reward approximation error analyzed in Proposition 6.1. Depending on how the “freeze error” and the “end-of-horizon error” interact, we see different behaviors of the overall reward approximation error, as illustrated in Figure 4. For both examples, the end-of-horizon error increases at first, but then decreases toward zero as $T$ becomes large. In the example with $L_{f}<1$ (left), the freeze error converges as $T$ increases; however, if $L_{f}>1$, the freeze error grows without bound. This leads to two distinct patterns in the overall reward approximation error: (1) when the freeze error is well-controlled, the total reward approximation error (blue) generally follows the unimodal shape of the end-of-horizon error (green); and (2) when the freeze error (orange) grows without bound, the reward approximation error (blue) follows a unimodal shape at first, but eventually grows again.
Regret of Base VI
We pause here briefly to state Proposition 7.1, a well-known property that gives the regret of exact VI on the base model. We use this result as a comparison in to the regret of FSVI in the next section.
Proposition 7.1.
we recall that $r_{\textnormal{max}}$ is an upper bound on the reward.
Proof.
See Appendix E.5. ∎
Regret versus Computational Cost for FSVI and VI: Selecting $T$
Next, we show how Proposition 7.1 (the regret of Base VI) and Theorem 6.1 (the regret of FSVI) can be combined to provide intuition about whether to use Base VI or FSVI (and if so, which $T$ and $k$?). We consider the trade-off between regret and computational cost, using the definition of cost given in Section 5.2. For FSVI, the concrete problem facing a practitioner is as follows: given a computational cost budget $B$, what combination of $(T,k)$ should be used? We make two observations: first, for a fixed $T$, regret decreases with $k$ (see Theorem 6.1), and second, for a fixed $T$, cost increases with $k$ (see Section 5.2). Therefore, for a fixed $T$, one should run FSVI for $k_{\text{max}}(T,B)$ iterations, where $k_{\text{max}}(T,B)$ is the largest $k$ such that the computational cost of running FSVI with $T$ is less than $B$. It thus follows that to select $T$, one should simply examine the regret for different values of $T$ at $k=k_{\text{max}}(T,B)$ and select the value of $T$ with the lowest regret.
The above procedure can be converted into a simple visualization. In Figure 5, we plot regret versus computational cost. This is done for both FSVI for different values of $T$ and also for standard value iteration (Base VI). Each colored curve traces the regret-cost trade-off generated by varying $k$ for a distinct $T$ (though $k$ is not shown). Imposing the computational budget corresponds to drawing the vertical line $x=B$ and where this line intersects a curve, one can read off the best attainable regret given budget $B$ for the algorithm represented by that curve. By scanning the intersection points and selecting the one with the lowest regret, one can identify the best algorithm to use, whether it is FSVI with a particular $T$ or simply Base VI. Hence, the plot allows us to visually solve the constrained optimization problem of minimizing regret under a given computational budget. This exercise leads us to interesting takeaways illustrated in Figure 5:
@@
Input: State distribution $\rho$, number of iterations $k$, number of state samples $N$, number of next-state samples $M$, and a regression procedure Regress.
@@ 3       Sample a dataset $\mathcal{D}=\{s^{(i)}\}_{i=1}^{N}\sim\rho$
5             $u^{(i)}=\max_{a}\,r(s^{(i)},a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k^{\prime}-1}\bigl{(}s^{(i,a,j)}\bigr{)}$, where $s^{(i,a,j)}$ is sampled from the generative model, conditional on state $s^{(i)}$ and action $a$.
@@ 7       $U^{k^{\prime}}=\texttt{Regress}(\{(s^{(i)},u^{(i)})\}_{i=1}^{N})$
10       $\nu^{k}(s)=\operatorname*{arg\,max}_{a}\,r(s,a)+\gamma\,\frac{1}{M}\sum_{j=1}^{M}\,U^{k}\bigl{(}s^{(a,j)}\bigr{)}$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$.
@@
Algorithm 3 Fitted Value Iteration (FVI) (Munos and Szepesvári, 2008)
Input: State distribution $\rho$, number of iterations $k$, number of state samples $N$, number of next-state samples $M_{l}$ (lower level), $M_{u}$ (upper level), and a regression procedure Regress.
@@
1 Initialize $\tilde{J}_{T}\equiv 0$.
2
3for $t=T-1,T-2,\ldots,1$ do
4       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
5       for each $(x^{(i)},y^{(i)})\in\mathcal{D}$ do
6             $J_{t}^{(i)}=\max_{a\in\mathcal{A}}\,r\bigl{(}x^{(i)},y^{(i)},a\bigr{)}+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}\tilde{J}_{t+1}\bigl{(}x^{(i)},y^{(i,a,j)}\bigr{)}$ where $y^{(i,a,j)}$ is sampled from the generative model, conditional on state $(x^{(i)},y^{(i)})$ and action $a$.
@@ 4       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
8       $\tilde{J}_{t}=\texttt{Regress}\Bigl{(}\bigl{\{}\bigl{(}(x^{(i)},y^{(i)}),J_{t}^{(i)}\bigr{)}\bigr{\}}_{i=1}^{N}\Bigr{)}$
@@
10Let $\tilde{\bm{\pi}}^{*}=(\pi_{1}^{*},\ldots\pi^{*}_{T-1})$, with $\pi_{t}^{*}(s)=\operatorname*{arg\,max}_{a}\,r(s,a)+\gamma\,\frac{1}{M_{l}}\sum_{j=1}^{M_{l}}\,\tilde{J}_{t+1}\bigl{(}s^{(a,j)}\bigr{)}$, where $s^{(a,j)}$ is sampled from the generative model, conditional on state $s$ and action $a$. Initialize $\hat{V}^{0}$ arbitrarily. for $k^{\prime}=1,2,\ldots,k$ do
11       Sample a dataset $\mathcal{D}=\{(x^{(i)},y^{(i)})\}_{i=1}^{N}\sim\rho$
@@
13             ${V}^{(k^{\prime})}=\max_{a\in\mathcal{A}}\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\left[\tilde{R}^{(i,a,j)}+\gamma^{T}{V}^{k^{\prime}-1}\left(x^{(i,a,j)},y^{(i,a,j)}\right)\right]$ where $x_{T}^{(i,a,j)}$, $y_{T}^{(i,a,j)}$, and $\tilde{R}^{(i,a,j)}$ are sampled using a $T$-step transition conditioned on $x^{(i)}$, $a$, $\tilde{\bm{\pi}}^{*}$ and $\tilde{J}_{1}$.
@@
17for $s_{0}=(x_{0},y_{0})\in\mathcal{S}$ do
18       $\tilde{\mu}^{k}(x_{0},y_{0})=\operatorname*{arg\,max}_{a\in\mathcal{A}}\frac{1}{M_{u}}\sum_{j=1}^{M_{u}}\left[\tilde{R}^{(i,a,j)}+\gamma^{T}\tilde{V}^{k}\left(x_{T}^{(i,a,j)},y_{T}^{(i,a,j)}\right)\right]$ where $x_{T}^{(i,a,j)}$, $y_{T}^{(i,a,j)}$, and $\tilde{R}^{(i,a,j)}$ are sampled from a $T$-step transition as above.
@@
So far, we have operated in the dynamic programming setting, where the model (rewards and transitions) is assumed to be known. We now move on to a more practical and realistic setting where we assume access to a generative model or simulator of the environment (also referred to as a restart distribution), an assumption frequently used in the RL literature (Kakade and Langford, 2002; Kearns et al., 2002; Munos and Szepesvári, 2008; Li et al., 2020; Agarwal et al., 2020). This setting is widely studied and captures many real-world use cases where it is possible to simulate environment transitions (Bellemare et al., 2012; Todorov et al., 2012; Brockman et al., 2016; Tassa et al., 2018), without having full access to closed-form model dynamics. This is an important setting that occupies a middle ground between dynamic programming and traditional model-free RL.
Fitted value iteration (FVI) is a simple method that approximates the value iteration backup step using regression (Munos and Szepesvári, 2008), which we keep as an abstract subroutine using the notation $\texttt{Regress}\bigl{(}\{(s^{(i)},y^{(i)})\}_{i=1}^{N}\bigr{)}$, which refers to the problem of regressing $\{y^{(i)}\}$ onto $\{s^{(i)}\}$. The Regress procedure is meant to encode all of the information needed to perform the regression: basis functions, approximation architecture (e.g., linear model or neural network), optimization procedure (e.g., stochastic gradient descent), etc. In Algorithm 3, we show the abstract version of FVI and in Algorithm 4, we give the frozen-state extension, Frozen-state FVI (FSFVI), which can be viewed as a practical extension of FSVI for the RL setting.
Additionally, in Appendix C, we present an alternative version of FVI that is less practical, but more amenable to theoretical analysis, leading to a regret bound that we provide in Theorem C.1. These results are motivated by the approximate value iteration method first presented in Tsitsiklis and Van Roy (1996). Given that both the method and analysis are technical and notation-heavy, we relegated the discussion to the Appendix.
@@
For our numerical experiments, we operate in the simulator-based RL setting described above, where a generative model or simulator is available but the MDP is not explicitly known. Note that both FVI and FSFVI naturally apply in this setting, as expectations are approximated via sample averages using either $M$, $M_{l}$, or $M_{u}$ next-state samples. In addition, we introduce empirical counterparts to Base VI (Algorithm 1) and FSVI (Algorithm 7), which we refer to as Base Empirical-VI (Base E-VI) and Empirical-FSVI (E-FSVI), following the naming convention of Haskell et al. (2016). These algorithms are identical to their exact counterparts, except that all expectations are replaced with sample averages. For completeness, we provide full specifications in Appendix A.
In Section 5.2, we analyzed the computational cost of FSVI under the assumption of a known model with exact expectation computations. In the RL setting, however, this definition requires adjustment. As a natural analogue, we use the number of value function evaluations as a proxy for computational effort. This makes sense in the RL context because value-based methods, whether value iteration, Q-iteration, or their fitted variants, all rely on repeatedly evaluating a value function (or Q-function). These evaluations typically dominate the overall computation time, especially when using function approximation. Thus, counting value function evaluations provides a measure of computational cost that aligns closely with RL practice, while being closely related to the exact results from Section 5.2.
We compare to a range of algorithms. For purely tabular methods, we consider:
  • E-FSVI with $T\in\{3,6,12\}$: Empirical-FSVI, as described in Algorithm 6. For each domain, we run E-FSVI for each $T$ value in order to illustrate the trade-offs.
  • Base E-VI: Empirical-VI, as described in Algorithm 5, applied to the “base”, non-hierarchical problem. We also use Base E-VI to compute an approximation of the optimal value.
  • Base E-QI: Empirical Q-iteration is the analogue of E-VI using a Q-function representation.
  • Slow-Agn Base E-VI: Slow-agnostic Base E-VI, a simple baseline that pretends the slow state does not exist and assumes the value function only depends on the fast state. [^8] To implement this method, we convert the “full” state to a “partial” state by deleting the slow state, and when needing to interact with the simulator, we convert from a “partial” state back to a “full” state by injecting a random value for the slow state. See Appendix B.1 for further details.
@@
  • {Lin, Deep}-FSFVI: Our proposed frozen-state fitted value iteration, as introduced in Algorithm 4, using either linear function approximator (Lin-FSFVI) or a three-layer neural network function approximator (Deep-FSVI).
  • {Lin, Deep}-FVI: Fitted value iteration (Munos and Szepesvári, 2008) using either linear function approximator (Lin-FVI) or a three-layer neural network function approximator (Deep-FVI). This algorithm is also known as least-square Monte Carlo, popularized by Longstaff and Schwartz (2001) and frequently used in the operations literature (see, e.g., Nadarajah et al. (2017)).
  • {Lin, Deep}-FQI: Fitted Q-iteration (Ernst et al., 2005) using either linear function approximator (Lin-FQI) or a three-layer neural network function approximator (Deep-FQI).
For Lin-FSFVI and Lin-FVI, we use the raw state feature vectors as input to a linear function approximator, without applying any feature transformation or basis functions. For Deep-FSFVI, FVI, we input the raw state vector into a fully connected feedforward neural network with two hidden layers, each comprising 4 units and ReLU activations, followed by a final linear output layer. For Lin-FQI and Deep-FQI, the input is formed by concatenating the raw state vector with a one-hot encoding of the discrete action (a standard practice in RL for handling discrete action spaces). The architecture of the Deep-FQI network is otherwise identical to that of Deep-FVI. All models are trained using the mean squared error loss and optimized with the Adam optimizer (Kingma, 2014). A detailed open-source implementation is also provided with this paper.
We test these algorithms across three domains, summarized in Table 2. Each domain uses a high discount factor of $\gamma=0.995$, corresponding to an effective planning horizon of approximately $2001/(1-\gamma)=$ 200 steps. We describe each experiment below; in each section, we use self-contained notation. Detailed experimental settings are given in Appendix B.
@@
Inventory control 561 11 6,171 Base E-VI, E-FSVI, Base E-QI, Slow-Agn Base E-VI
Spatial-task gridworld 4,356 32 139,392 Base E-VI, E-FSVI, Base E-QI, Slow-Agn Base E-VI
@@
where $d_{t+1}$ is drawn from a distribution $P_{d}$ conditional on $d_{t}$. The (expected) reward in each period is given by:
@@
which represents the revenue, variable order cost, and fixed order cost. The objective is to maximize infinite horizon discounted reward, with $\gamma=0.995$. In our experiment, we set $y_{\textnormal{max}}=50$, $\mathcal{A}_{\textnormal{order}}=\{0,5,\ldots,50\}$, $d_{t+1}=\textnormal{clip}_{[0,50]}(d_{t}+\epsilon_{t+1})$, and $\{\epsilon_{t}\}$ are i.i.d. random variables that take value 0 with probability $0.8$ and ±$\pm 1$ with probability 0.1 each. [^9] We define $\textnormal{clip}_{[a,b]}(x)=\max\{\min\{x,b\},a\}$. We use $M=M_{u}=50$ for all methods, and $M_{l}=1$ for E-FSVI because the lower-level problem is deterministic.
Figure 6: Mean return (with 95% CI) versus computational cost for the inventory problem across five independent replications. The E-FSVI, we plot the performance of the policy after each lower-level value iteration and 3 times per upper-level value iteration. For the other methods, we plot 3 times per value iteration. To compute the optimal value, we ran Base E-VI for 50 iterations with $M=100$ samples per backup.
The results, shown in Figure 6, demonstrate that frozen-state methods substantially outperform standard approaches in terms of computational efficiency. Specifically, E-FSVI with moderate values of $T$, such as $T=6$ and $T=12$, achieves near-optimal return with significantly fewer value function evaluations compared to Base E-VI and E-QI. This supports the intuition that freezing the slow-moving demand state $d_{t}$ over short planning horizons can preserve essential structure while simplifying the lower-level planning problem.
Among the E-FSVI variants, we find that $T=6$ offers the best overall performance when considering both reward and computational cost. Larger values of $T$, such as $T=12$, initially converge faster but eventually suffer from increased approximation error due to longer freezing horizons. Conversely, if we look at the earlier part of the curves, E-FSVI with $T=3$ seems to benefit less from the improved discount factor.
While Base E-VI and E-QI eventually reach optimal performance, they require significantly more computation. In contrast, the slow-agnostic baseline performs poorly, confirming that ignoring the slow state leads to suboptimal policies. These findings reinforce both our theoretical claims from Section 6 and the trade-offs we observed in Figure 5 from Section 7.